# Ilias Bilionis

### v1

@@ 7 Conclusions
This study introduced a Bayesian calibration framework aimed at improving the transferability of engine-out NOx predictive models across different engines. The calibrated predictions obtained through this methodology demonstrated substantial accuracy improvements compared to the baseline GP model, consistently observed across multiple engines and diverse operating cycles. Specifically, posterior predictive medians exhibited closer alignment with measured NOx emissions, and their associated credible intervals effectively captured the uncertainty due to engine-to-engine variability. The cumulative NOx analysis and reduction in percentile errors further substantiated the robustness and practical value of the proposed approach for real world applications demanding cross engine generalizability.
@@
  • The current model assumes that the sensor and output biases ($\alpha_{i},b_{i}$) are time invariant (constant offsets) over the entire operating cycle. In reality, biases might exhibit time dependent drift or vary non linearly with operating conditions, which is not captured by this framework.
@@
Shrenik Zinage: Methodology, Software, Validation, Visualization, Writing - original draft. Peter Meckl: Funding acquisition, Writing - review and editing. Ilias Bilionis: Funding acquisition, Methodology, Validation, Writing - review and editing.

### v2

@@ 7 Conclusions
This study introduced a Bayesian calibration framework aimed at improving the transferability of engine-out NOx predictive models across different engines. The calibrated predictions obtained through this methodology demonstrated substantial accuracy improvements compared to the baseline GP model, consistently observed across multiple engines and diverse operating cycles. Specifically, posterior predictive medians exhibited closer alignment with measured engine-out NOx, and their associated credible intervals effectively captured the uncertainty due to engine-to-engine variability. The cumulative NOx analysis and reduction in percentile errors further substantiated the robustness and practical value of the proposed approach for real world applications demanding cross engine generalizability.
@@
  • The current model assumes that the sensor and output biases ($\alpha_{i},b_{i}$) are time invariant (constant offsets) over the entire operating cycle. While empirical evidence in Figs. 3 and 4 supports this approximation over the analyzed windows, real-world deployment over extended durations may introduce time-varying bias due to sensor aging, thermal cycling, fouling, or actuator wear. In such cases, a static calibration may gradually lose accuracy as drift accumulates. The primary sources of error would arise from unmodeled slow bias drift rather than GP model error. Depending on drift magnitude, this could manifest as gradual increases in RMSE and systematic deviations in cumulative NOx over long operating horizons.
Potential extensions of this framework could address this by incorporating time dependent bias terms, potentially modeled via hierarchical GP or by implementing an online calibration strategy that periodically updates the bias estimates $(\alpha_{i},b_{i})$ using a sliding window of recent operating data.
Note that although this study focuses on engine out NOx, the proposed framework is more broadly applicable to systems in which a pretrained surrogate is deployed across a heterogeneous population and suffers from sensor specific biases. Examples include predictive models for other emission species such as particulate matter, carbon monoxide, or unburned hydrocarbons, as well as models for combustion phasing, turbocharger speed, or aftertreatment state estimation.
@@
Shrenik Zinage: Methodology, Software, Validation, Visualization, Writing - original draft. Peter Meckl: Funding acquisition, Writing - review and editing. Ilias Bilionis: Funding acquisition, Methodology, Writing - review and editing.